# A Review on Single-Problem Multi-Attempt Heuristic Optimization

### v1

@@ 6. Framework utility and design of new strategies
The tables provide easy access to selection strategies associated with different components. For example, a reader interested in multi-start (MS) strategies can identify relevant works in Table 3 by checking which rows are filled in the $p_{\Xi}$ column. The corresponding section for each strategy can then be located using the last columns of Table 3, Table 5, and Table 6, which indicate where each strategy is discussed. Based on this information, the reader can gain a general understanding of the main characteristics of each strategy and choose the one that best aligns with their specific requirements and preferences. Then, the reader may identify several MS strategies suitable for their problem, such as randomized restarts (Cicirello and Smith, 2005), Bayesian optimization-based restarts (Mathesen et al., 2021), or kernel density estimation-based strategies (Luersen and Le Riche, 2004), and gain a detailed understanding of each approach by consulting the corresponding references in the literature. Moreover, since an alternative is defined as any combination of the four components, some readers may also be interested in strategies that address the selection of multiple components simultaneously. For instance, a reader seeking both parameter tuning (PT) and multi-start (MS) may combine a configuration selection strategy from one work with an initialization selection strategy from another. Alternatively, the reader may adapt strategies originally developed for one component to reuse them in the selection of a different component of interest. Indeed, we observe that the same strategies are often applied to different components across various works. For example, Bayesian optimization is employed for both configuration (Klein et al., 2017; Hu et al., 2022) and initialization (Mathesen et al., 2021; Meng et al., 2023), while exponential-increasing strategies are associated with both configuration (Auger and Hansen, 2005; Doerr and Zheng, 2020a) and stopping criteria (Huang et al., 2020b; Li et al., 2018).
@@ 7. Conclusion
Although SIMHO is commonly encountered in practice, to the best of our knowledge, this is the first review that focuses exclusively on this problem and unifies alternative selection strategies that address the selection across all four components. Selection strategies suitable for SIMHO have been studied in isolation within the topics of algorithm selection, parameter tuning, multi-start, and resource allocation. Existing reviews fail to provide a comprehensive view, either because they do not consider strategies from all four topics or because they include others that are not suitable for SIMHO scenario. As a result, practitioners aiming to address SIMHO face limitations in understanding and applying the broad range of existing strategies. This review addresses that gap by bringing together strategies that, although developed independently, are all suitable for tackling SIMHO. We formalize these strategies within a unified framework using consistent terminology, and introduce a taxonomy that organizes and characterizes them accordingly. This framework-based review thus facilitates both the identification and the development of selection strategies for SIMHO.

### v2

@@ 6.1. Framework utility and design of new strategies
The tables provide easy access to selection strategies associated with different components. For example, a reader interested in multi-start (MS) strategies can identify relevant works in Table 3 by checking which rows are filled in the $p_{\Xi}$ column. The corresponding section for each adaptive strategy can then be located using the last columns of Table 3, Table 5, and Table 6, which indicate where each strategy is discussed. Specifically, papers that have a checkmark in EMQ or EQI of the last column of Table 3 are described in one of the items in Sections 5.3.1 or 5.3.2, respectively, and this item is completely determined by the checkmarks in the last columns of Table 5 and Table 6 associated with that paper. Based on this information, the reader can gain a general understanding of the main characteristics of each strategy and choose the one that best aligns with their specific requirements and preferences. Then, the reader may identify several MS strategies suitable for their problem, such as randomized restarts (Cicirello and Smith, 2005), Bayesian optimization-based restarts (Mathesen et al., 2021), or kernel density estimation-based strategies (Luersen and Le Riche, 2004), and gain a detailed understanding of each approach by consulting the corresponding references in the literature. Moreover, since an alternative is defined as any combination of the four components, some readers may also be interested in strategies that address the selection of multiple components simultaneously. For instance, a reader seeking both parameter tuning (PT) and multi-start (MS) may combine a configuration selection strategy from one work with an initialization selection strategy from another. Alternatively, the reader may adapt strategies originally developed for one component to reuse them in the selection of a different component of interest. Indeed, we observe that the same strategies are often applied to different components across various works. For example, Bayesian optimization is employed for both configuration (Klein et al., 2017; Hu et al., 2022) and initialization (Mathesen et al., 2021; Meng et al., 2023), while exponential-increasing strategies are associated with both configuration (Auger and Hansen, 2005; Doerr and Zheng, 2020a) and stopping criteria (Huang et al., 2020b; Li et al., 2018).
@@
6.2. Transferability of strategies across settings and scope of the review
From a purely formal perspective, one could argue that the single-problem setting may be seen as a particular variant of the multi-problem setting in which only one instance is available (see Appendix B for a more detailed discussion). However, this formal inclusion does not imply methodological equivalence or transferability.
Strategies designed for multi-problem settings are inherently built around the availability of multiple instances and the possibility of transfer or generalization across them. When these assumptions do not hold, as in the single-problem setting, such strategies either cannot be applied effectively or change their objective and interpretation. For example, per-instance algorithm selection (Rice, 1976) relies on supervised models trained on features extracted from a set of problem instances. When only a single problem is available, such models cannot be constructed efficiently due to the lack of training instances or sufficiently related problems. In contrast, methods such as irace (López-Ibáñez et al., 2016), which are designed to identify configurations that perform robustly across multiple instances, can be adapted to the single-problem setting by repeatedly racing configurations on the same problem (Bischl et al., 2023). In this case, the strategy effectively shifts into a resource allocation or successive halving scheme, rather than a method for identifying configurations that are robust across instances. This is conceptually similar to approaches such as Hyperband (Li et al., 2018), which are included in our review.
Whether the single-problem setting should be considered a niche therefore strongly depends on the target audience. For practitioners and researchers interested in single-problem optimization, existing reviews leave a clear gap: they either study strategies originally designed for this setting separately or include multi-problem strategies that are not efficiently transferable to the single-problem case. By unifying and systematizing the strategies that were originally designed for the single-problem setting, our review directly addresses this gap and provides a focused and coherent reference for this audience.
Conversely, our review is not intended for users interested in multi-problem optimization, as it is intentionally and exclusively focused on the single-problem setting. Moreover, single-problem strategies are generally not efficiently transferable to the multi-problem case either. For instance, such strategies could be applied in a plug-and-play manner to each instance independently, but this would ignore cross-instance information and therefore fail to exploit robustness or transfer. Similarly, in multi-start strategies (Luersen and Le Riche, 2004; Mathesen et al., 2021) that progressively learn the most interesting regions of the search space to start a local search, changing the problem would not be meaningful, because the learned information that is used to guide new local search runs is specific to the problem instance and its search space.
Overall, this review is intentionally and exclusively focused on the single-problem setting, motivated by its practical relevance and by the lack of a unified and systematic review dedicated to it. As a consequence of this deliberate focus, the review does not aim to address the needs of practitioners interested in multi-problem optimization. That setting is already covered by existing surveys, some of which are listed in Table 1, although evaluating or recommending them falls outside the scope of this work.
@@ 7. Conclusion
Although SIMHO is commonly encountered in practice, to the best of our knowledge, this is the first review that focuses exclusively on this problem and unifies alternative selection strategies that address the selection across all four components. Selection strategies suitable for SIMHO have been studied separately within the topics of algorithm selection, parameter tuning, multi-start, and resource allocation. Existing reviews fail to provide a comprehensive view, either because they do not consider strategies from all four topics or because they include others that are not suitable for SIMHO scenario. As a result, practitioners aiming to address SIMHO face limitations in understanding and applying the broad range of existing strategies. This review addresses that gap by bringing together strategies that, although developed independently, are all suitable for tackling SIMHO. We formalize these strategies within a unified framework using consistent terminology, and introduce a taxonomy that organizes and characterizes them accordingly. This framework-based review thus facilitates both the identification and the development of selection strategies for SIMHO.
@@ Appendix A Hierarchical dependence between components
Appendix B Formal similarity between single- and multi-problem settings
Problem (4) provides a unified formalization of the optimization problem underlying the single-problem multi-attempt setting considered in this review. Since this paper is a review focused entirely on the single-problem setting, the main role of the formalization, in Section 3, is to establish a shared terminology that allows us to systematically describe, in Section 5, selection strategies that were originally studied separately but are all suitable for this setting. Therefore, this formalization, which serves as a supporting tool to make this review more structured and coherent, is defined specifically for the single-problem setting and does not directly extend to the multi-problem setting, despite possible formal similarities that may be observed. It is important to emphasize that the framework is also not intended to model the multi-problem setting, since from the outset the main motivation of this review is the exclusive focus on SIMHO due to its practical relevance, as described in Section 1.
Specifically, in our framework, Problem (4) formalizes the optimization objective of identifying the alternative $a\in A$ that yields the best performance for a single optimization problem defined by a given objective function $f$ and search space $X$, i.e., an objective of the form $\operatorname*{\arg\max}_{a\in A}~f_{a}$. In contrast, optimization problems in the multi-problem setting involve fundamentally different objectives that are defined over collections of problem instances rather than a single one, as discussed in Section 2. For example, in per-instance algorithm selection (Rice, 1976), the objective is to identify, for each problem instance, the algorithm that performs best. The learned mapping from instances to algorithms is the strategy used to achieve this goal, but the underlying objective would correspond to an extension of Problem (4) over a set of objective functions (i.e., a “for all” over multiple $f$’s), which is not captured by our formulation. Similarly, methods such as irace (López-Ibáñez et al., 2016) aim to identify configurations that perform robustly across a collection of instances, which would require a multi-objective or aggregated optimization objective incorporating performance over all instances of interest. In both cases, the availability of multiple problem instances is an explicit assumption, which does not align with the Problem (1) underlying Problem (4).